# Measuring Self-Rating Bias in LLM-Generated Survey Data: A Semantic Similarity Framework for Independent Scale Mapping

### v1

@@ 4.8.2 LLM Baseline Comparison
A notable divergence emerges in label sensitivity. Claude exhibits complete label invariance: both variants produce identical global accuracy (87%), though they differ on which specific cases they miss. GPT-4o, in contrast, shows a 3-percentage-point accuracy decrease with domain-aware labels (83% $\to$ 80%), suggesting that question-specific scale anchors introduce slight miscalibration in GPT-4o’s scale mapping. This is the opposite of what one might expect—more specific labels should help—and suggests that GPT-4o’s internal scale representation is well-calibrated for generic evaluative language but slightly disrupted by domain-specific terminology. This finding nuances the label invariance observation from Section 4.7: label invariance holds for Claude but may not generalize to all model families.
@@ 4.9.1 Motivation
The baseline comparison (Section 4.7) established that direct LLM rating achieves higher accuracy than SSR on human-written text. However, that comparison does not test the circularity scenario that motivates SSR: a model rating its own generated text. When the same model generates and evaluates a response, the evaluation may reflect the model’s internal consistency—its tendency to produce text that aligns with its own classification boundaries—rather than the text’s true semantic content. This experiment directly tests whether self-rating produces measurable bias.
@@ 4.9.2 Design
We generated synthetic survey responses for all 69 reference test cases using Claude Haiku 4.5 with LLM generation temperature $T=0.7$ (distinct from the SSR softmax temperature $\tau$) and five persona profiles (Table 17) varying in communication style. Each persona generated responses at all target rating levels (1–5), producing $69\times 5=345$ generated texts. Each text was then independently rated by two methods:
@@
Table 17: Persona profiles used for text generation. Each persona generated one response per test case ($69\times 5=345$ total texts).
Persona Style Description
Maria Detailed, analytical A 35-year-old marketing manager who writes thorough, well-structured responses with specific details
James Blunt, direct A 58-year-old retired engineer who says exactly what he thinks without softening
Alex Hedging, qualified A 24-year-old graduate student who qualifies and hedges opinions
Linda Practical, concise A 42-year-old nurse and mother of two who gives short, practical answers
Ryan Informal, expressive A 21-year-old barista and college student who writes casually and expressively
@@
Both methods show comparable accuracy degradation when moving from human to generated text (14-14 pp for LLM, 15-15 pp for SSR). The near-identical magnitude of this degradation ($\Delta\Delta=1$ pp) suggests that generated text is inherently harder to rate regardless of method—the difficulty is a property of the text, not the rating approach. The LLM’s accuracy advantage over SSR is stable across text sources (+22 pp on human text, +23 pp on generated text), indicating no interaction between text source and rating method at the level of aggregate accuracy. However, the nature of the errors differs fundamentally, as the confirmatory tests below demonstrate.
@@ 4.9.3 Results
LLM self-rating shows a consistent downward bias relative to reference ratings (mean signed error =0.23=-0.23, $\text{SD}=0.46$), while SSR shows near-zero bias (mean signed error =+0.08=+0.08, $\text{SD}=0.93$). This pattern—the LLM systematically underestimating the intended rating when evaluating its own generated text—is the opposite of the naïve prediction that self-rating would produce inflated agreement. We hypothesize that this reflects a temperature asymmetry between generation and evaluation: the generator ($T=0.7$) samples from a broader distribution that includes more expressive, emotionally vivid language, while the rater ($T=0$) deterministically selects the most conservative interpretation of that text. In effect, the generator “writes like a 4” but the rater “reads like a 3.” This is a testable hypothesis: if correct, increasing the rater’s temperature should reduce the negative bias (at the cost of increased variance), and decreasing the generator’s temperature should produce text that the rater scores more accurately. Testing this hypothesis is a priority for future work but is beyond the scope of the present study.
@@
LLM rating errors show 4-fold variance compression: $\sigma^{2}=0.21$ compared to SSR’s $\sigma^{2}=0.87$ (Levene $F(1,688)=60.0$, $p<.001$). This means LLM errors cluster tightly around a small negative offset, while SSR errors are more dispersed but centered near zero. Crucially, the cross-model control (Section 4.9.5) shows that GPT-4o produces nearly identical compression when rating Claude-generated text ($\sigma^{2}=0.23$, within/cross ratio =0.93=0.93), indicating that this effect is a general property of LLM-based rating rather than a consequence of a model evaluating its own output. The compression likely reflects how autoregressive models process evaluative text—converging on similar scale interpretations regardless of the text’s origin—rather than shared internal representations between generator and rater.
@@ 4.9.4 Exploratory Analyses
Table 20 shows per-domain accuracy and divergence on generated text. Two domains exhibit a reversal where SSR outperforms LLM self-rating: likelihood (57% vs. 53%) and purchase intent (70% vs. 57%). Notably, purchase intent is the construct for which maier2024ssr demonstrated distributional validity against real human data, and it is also a domain where SSR achieves strong performance on human text (88%, Section 4.6).
@@ 5.1 Key Findings
5.1.1 Circularity Can Be Substantially Reduced—At a Cost
@@
The appropriate interpretation depends on the research context. When the goal is maximizing accuracy, direct LLM rating is more accurate. When the goal is methodological independence, SSR provides a guarantee that no amount of LLM accuracy can substitute. These are complementary, not competing, objectives. The cross-model comparison (Section 4.8) confirms this tradeoff is robust across model families, and the circularity experiment with cross-model control (Section 4.9) demonstrates that LLM-based rating produces systematically different error patterns than independent measurement—not just within-model, but across model families.
@@ 5.1.2 Anchor Quality Dominates All Other Factors
The dominance of anchor quality over all algorithmic manipulations connects directly to the survey methodology literature on behavioral anchoring (krosnick2010question; tourangeau2000psychology): just as human respondents are more reliable with concrete behavioral scale anchors, embedding models produce more discriminative vectors when anchors contain distinctive, specific language. The mechanism is analogous—richer semantic content provides better “targets” for mapping, whether the mapping is cognitive (human) or geometric (embedding). This principle extends beyond SSR to any embedding-based text classification against semantic prototypes.
@@ 5.1.3 Mapping Error Magnitude in Context
The accuracy gap reflects a fundamental difference in rating mechanism: autoregressive models have internalized ordinal mapping from pretraining, while embedding models must reconstruct it geometrically. The information asymmetry control (Table 14) rules out the hypothesis that LLMs simply have access to more information. Crucially, the gap is concentrated in weaker domains; for well-anchored domains (satisfaction, purchase intent), all methods achieve $\geq$80% exact. The 12 pp improvement from Voyage to OpenAI embeddings—without any change to the framework—demonstrates that model selection is a high-leverage path that does not compromise independence.
For context, human test-retest reliability on 5-point Likert scales typically produces response standard deviations of 0.5–1.0 (nunnally1994psychometric; krosnick2010question), though test-retest measures intra-rater stability while our evaluation measures inter-method agreement (see Limitations).
@@
Our purchase intent results (88% exact, MAE =0.13=0.13) converge with maier2024ssr’s (maier2024ssr) distributional validation (KS similarity >0.85>0.85) for the same construct, providing indirect cross-study evidence that SSR generalizes across research groups, embedding models, and evaluation methodologies. However, individual-level accuracy and distributional validity measure fundamentally different properties: a method with 65% individual exact match can still produce excellent distributional fit if errors are symmetric, while a method with high individual accuracy but systematic bias would fail distributional validation. Bridging these two levels—case-level precision and population-level fidelity—is a priority for future work.
@@ 5.1.4 Label-Invariant LLM Accuracy
Label invariance—identical LLM accuracy regardless of scale label formulation—deepens the circularity concern: if LLMs have internalized ordinal mapping as a general capability, then text generation and scale assignment draw on the same learned construct. The cross-model comparison reveals this is not universal (Claude: 87% invariant; GPT-4o: 3 pp decrease with domain-aware labels), suggesting that the severity of internalized psychometric knowledge varies across model families. A practical implication is that the rating step does not require expensive models: Claude Haiku 4.5 reaches the estimated arguability ceiling ($\sim$87–90%; Section 5.2.7).
@@ 5.2.1 Single-Rater Ground Truth
This is the most significant limitation of the present work. All expected ratings were assigned by a single researcher, not derived from multi-rater coding or crowd-sourced consensus. Typical inter-rater reliability for Likert-scale coding ranges from $\kappa=0.60$$0.80$ (landis1977measurement), and for subtle cases (e.g., mixed-signal responses near the boundary between ratings 2 and 3), reasonable coders might disagree on the “correct” rating. Our reported accuracy metrics therefore conflate SSR error with the uncertainty in the ground truth itself. A multi-rater validation study with established inter-rater agreement (e.g., Krippendorff’s $\alpha\geq 0.80$) is planned for the journal version of this paper and would provide more robust benchmarks for SSR’s true accuracy.
@@ 5.4 Future Work
    Distributional validation against human data: maier2024ssr demonstrated that SSR achieves 90% of human test–retest reliability at the population level, but their evaluation and ours use different embedding models, anchor strategies, and evaluation metrics. The critical next step is to apply our calibrated SSR configuration (naturalistic anchors, asymmetric embedding) to real survey instruments with human response data, enabling direct distributional comparison (KS tests, distribution shape analysis) alongside the individual-level metrics reported here. This would bridge the case-level precision established in the present work with the population-level fidelity demonstrated by maier2024ssr, providing a complete validation picture.
@@ 6 Conclusion
These findings indicate that measurement independence is not merely desirable but practically necessary for accurate uncertainty quantification in synthetic survey data. The 12 pp accuracy improvement from Voyage to OpenAI embeddings—without changing the mathematical framework—demonstrates that model selection is a high-leverage path for closing the accuracy gap while preserving independence.
@@ Acknowledgments
Appendix A Anchor Library Summary
Table LABEL:tab:anchor-summary presents the extreme anchor statements (ratings 1 and 5) for each of the 15 semantic families in the SSR framework. All anchors are written in first-person naturalistic behavioral language, following the design principles described in Section 4.3. The complete library (15 families $\times$ 5 ratings =75=75 statements) is available in the project repository (see Data Availability).
Table 22: Anchor library summary: extreme anchors (ratings 1 and 5) for each semantic family. The full 75-statement library (including ratings 2–4) is available in the project repository.
Family Rating 1 (Lowest) Rating 5 (Highest)
Agreement Completely wrong. I strongly disagree, this doesn’t match reality at all… Absolutely right. I completely agree, this is exactly correct…
Satisfaction Terrible. I was frustrated and angry the entire time, nothing worked right… Excellent. Everything was smooth, fast, and easy. I’m really impressed…
Likelihood No, I definitely won’t. I had a bad experience and I have no intention… Absolutely, I definitely will. I had a great experience and I’m already looking forward…
Ease Incredibly frustrating. I kept getting stuck and couldn’t figure out what to do… Couldn’t have been simpler. Everything was obvious and intuitive…
Importance Couldn’t care less about this. It has zero impact on my life… This is absolutely essential. I can’t imagine doing without it…
Familiarity Never heard of it. This is completely new to me… I know this inside and out. I use it all the time…
Appeal Hard pass. Nothing about this interests me at all… I absolutely love this. I’m excited just thinking about it…
Value Total rip-off. I feel cheated paying this much for what you actually get… Incredible deal. I can’t believe how much you get for this price…
Trust I don’t trust this one bit. Something feels off… I trust this completely. It has a solid track record…
Quality Terrible quality. It feels cheap, flimsy, and poorly made… Outstanding quality. Premium feel in every way, this is clearly best-in-class…
Frequency Never, not once. This just isn’t something that’s part of my routine… All the time, practically every day. I can’t imagine not doing this…
Uniqueness Nothing special about this at all. It’s basically a copy of what’s already out there… Truly one of a kind. I’ve never seen anything quite like this before…
Relevance This has nothing to do with me. It’s completely off-base… Spot on, exactly what I was looking for. This fits my needs perfectly…
Impression Awful first impression. Everything about it rubs me the wrong way… Absolutely fantastic impression. I’m blown away by everything I’ve seen…
Purchase Intent No chance I’m buying this. It’s not for me at all… I’m buying this for sure. Take my money right now…

### v2

@@ 4.8.2 LLM Baseline Comparison
A notable divergence emerges in label sensitivity. Claude exhibits complete label invariance: both variants produce identical global accuracy (87%), though they differ on which specific cases they miss. GPT-4o, in contrast, shows a 3-percentage-point accuracy decrease with domain-aware labels (83% $\to$ 80%), suggesting that question-specific scale anchors introduce slight miscalibration in GPT-4o’s scale mapping. This is the opposite of what one might expect (more specific labels should help) and suggests that GPT-4o’s internal scale representation is well-calibrated for generic evaluative language but slightly disrupted by domain-specific terminology. This finding nuances the label invariance observation from Section 4.7: label invariance holds for Claude but may not generalize to all model families.
@@ 4.9.1 Motivation
The baseline comparison (Section 4.7) established that direct LLM rating achieves higher accuracy than SSR on human-written text. However, that comparison does not test the circularity scenario that motivates SSR: a model rating its own generated text. When the same model generates and evaluates a response, the evaluation may reflect the model’s internal consistency, its tendency to produce text that aligns with its own classification boundaries, rather than the text’s true semantic content. This experiment directly tests whether self-rating produces measurable bias.
@@ 4.9.2 Design
We generated synthetic survey responses for all 69 reference test cases using Claude Haiku 4.5 with LLM generation temperature $T=0.7$ (distinct from the SSR softmax temperature $\tau$) and five persona profiles varying in communication style (detailed/analytical, blunt/direct, hedging/qualified, practical/concise, and informal/expressive). Each persona generated responses at all target rating levels (1–5), producing $69\times 5=345$ generated texts. Each text was then independently rated by two methods:
@@
Both methods show comparable accuracy degradation when moving from human to generated text (14-14 pp for LLM, 15-15 pp for SSR). The near-identical magnitude of this degradation ($\Delta\Delta=1$ pp) suggests that generated text is inherently harder to rate regardless of method; the difficulty is a property of the text, not the rating approach. The LLM’s accuracy advantage over SSR is stable across text sources (+22 pp on human text, +23 pp on generated text), indicating no interaction between text source and rating method at the level of aggregate accuracy. However, the nature of the errors differs fundamentally, as the confirmatory tests below demonstrate.
@@ 4.9.3 Results
LLM self-rating shows a consistent downward bias relative to reference ratings (mean signed error =0.23=-0.23, $\text{SD}=0.46$), while SSR shows near-zero bias (mean signed error =+0.08=+0.08, $\text{SD}=0.93$). This pattern, the LLM systematically underestimating the intended rating when evaluating its own generated text, is the opposite of the naïve prediction that self-rating would produce inflated agreement. We hypothesize that this reflects a temperature asymmetry between generation and evaluation: the generator ($T=0.7$) samples from a broader distribution that includes more expressive, emotionally vivid language, while the rater ($T=0$) deterministically selects the most conservative interpretation of that text. In effect, the generator “writes like a 4” but the rater “reads like a 3.” This is a testable hypothesis: if correct, increasing the rater’s temperature should reduce the negative bias (at the cost of increased variance), and decreasing the generator’s temperature should produce text that the rater scores more accurately. Testing this hypothesis is a priority for future work but is beyond the scope of the present study.
@@
LLM rating errors show 4-fold variance compression: $\sigma^{2}=0.21$ compared to SSR’s $\sigma^{2}=0.87$ (Levene $F(1,688)=60.0$, $p<.001$). This means LLM errors cluster tightly around a small negative offset, while SSR errors are more dispersed but centered near zero. Crucially, the cross-model control (Section 4.9.5) shows that GPT-4o produces nearly identical compression when rating Claude-generated text ($\sigma^{2}=0.23$, within/cross ratio =0.93=0.93), indicating that this effect is a general property of LLM-based rating rather than a consequence of a model evaluating its own output. The compression likely reflects how autoregressive models process evaluative text, converging on similar scale interpretations regardless of the text’s origin, rather than shared internal representations between generator and rater.
@@ 4.9.4 Exploratory Analyses
Table 19 shows per-domain accuracy and divergence on generated text. Two domains exhibit a reversal where SSR outperforms LLM self-rating: likelihood (57% vs. 53%) and purchase intent (70% vs. 57%). Notably, purchase intent is the construct for which Maier et al., (2024) demonstrated distributional validity against real human data, and it is also a domain where SSR achieves strong performance on human text (88%, Section 4.6).
@@ 5.1 Key Findings
5.1.1 Circularity Can Be Substantially Reduced: At a Cost
@@
The appropriate interpretation depends on the research context. When the goal is maximizing accuracy, direct LLM rating is more accurate. When the goal is methodological independence, SSR provides a guarantee that no amount of LLM accuracy can substitute. These are complementary, not competing, objectives. The cross-model comparison (Section 4.8) confirms this tradeoff is robust across model families, and the circularity experiment with cross-model control (Section 4.9) demonstrates that LLM-based rating produces systematically different error patterns than independent measurement, not just within-model, but across model families.
@@ 5.1.2 Anchor Quality Dominates All Other Factors
The dominance of anchor quality over all algorithmic manipulations connects directly to the survey methodology literature on behavioral anchoring (Krosnick and Presser,, 2010; Tourangeau et al.,, 2000): just as human respondents are more reliable with concrete behavioral scale anchors, embedding models produce more discriminative vectors when anchors contain distinctive, specific language. The mechanism is analogous: richer semantic content provides better “targets” for mapping, whether the mapping is cognitive (human) or geometric (embedding). This principle extends beyond SSR to any embedding-based text classification against semantic prototypes.
@@ 5.1.3 Mapping Error Magnitude in Context
The accuracy gap reflects a fundamental difference in rating mechanism: autoregressive models have internalized ordinal mapping from pretraining, while embedding models must reconstruct it geometrically. The information asymmetry control (Table 14) rules out the hypothesis that LLMs simply have access to more information. Crucially, the gap is concentrated in weaker domains; for well-anchored domains (satisfaction, purchase intent), all methods achieve $\geq$80% exact. The 12 pp improvement from Voyage to OpenAI embeddings, without any change to the framework, demonstrates that model selection is a high-leverage path that does not compromise independence.
For context, human test-retest reliability on 5-point Likert scales typically produces response standard deviations of 0.5–1.0 (Nunnally and Bernstein,, 1994; Krosnick and Presser,, 2010), though test-retest measures intra-rater stability while our evaluation measures inter-method agreement (see Limitations).
@@
Our purchase intent results (88% exact, MAE =0.13=0.13) converge with Maier et al.,’s (2024) distributional validation (KS similarity >0.85>0.85) for the same construct, providing indirect cross-study evidence that SSR generalizes across research groups, embedding models, and evaluation methodologies. However, individual-level accuracy and distributional validity measure fundamentally different properties: a method with 65% individual exact match can still produce excellent distributional fit if errors are symmetric, while a method with high individual accuracy but systematic bias would fail distributional validation. Bridging these two levels, case-level precision and population-level fidelity, is a priority for future work.
@@ 5.1.4 Label-Invariant LLM Accuracy
Label invariance, identical LLM accuracy regardless of scale label formulation, deepens the circularity concern: if LLMs have internalized ordinal mapping as a general capability, then text generation and scale assignment draw on the same learned construct. The cross-model comparison reveals this is not universal (Claude: 87% invariant; GPT-4o: 3 pp decrease with domain-aware labels), suggesting that the severity of internalized psychometric knowledge varies across model families. A practical implication is that the rating step does not require expensive models: Claude Haiku 4.5 reaches the estimated arguability ceiling ($\sim$87–90%; Section 5.2.7).
@@ 5.2.1 Single-Rater Ground Truth
This is the most significant limitation of the present work. All expected ratings were assigned by a single researcher, not derived from multi-rater coding or crowd-sourced consensus. Typical inter-rater reliability for Likert-scale coding ranges from $\kappa=0.60$$0.80$ (Landis and Koch,, 1977), and for subtle cases (e.g., mixed-signal responses near the boundary between ratings 2 and 3), reasonable coders might disagree on the “correct” rating. Our reported accuracy metrics therefore conflate SSR error with the uncertainty in the ground truth itself. A multi-rater validation study with established inter-rater agreement (e.g., Krippendorff’s $\alpha\geq 0.80$) is planned for the journal version of this paper and would provide more robust benchmarks for SSR’s true accuracy.
@@ 5.4 Future Work
    Distributional validation against human data: Maier et al., (2024) demonstrated that SSR achieves 90% of human test–retest reliability at the population level, but their evaluation and ours use different embedding models, anchor strategies, and evaluation metrics. The critical next step is to apply our calibrated SSR configuration (naturalistic anchors, asymmetric embedding) to real survey instruments with human response data, enabling direct distributional comparison (KS tests, distribution shape analysis) alongside the individual-level metrics reported here. This would bridge the case-level precision established in the present work with the population-level fidelity demonstrated by Maier et al.,, providing a complete validation picture.
@@ 6 Conclusion
These findings indicate that measurement independence is not merely desirable but practically necessary for accurate uncertainty quantification in synthetic survey data. The 12 pp accuracy improvement from Voyage to OpenAI embeddings, without changing the mathematical framework, demonstrates that model selection is a high-leverage path for closing the accuracy gap while preserving independence.